OpenStudy (anonymous):

WILL FAN AND MEDAL! What is the slope of the line with the equation 2x-4y=4?

Vocaloid (vocaloid):

divide the entire equation by 4 first

OpenStudy (anonymous):

two cant be divided by four @vocaloid

Vocaloid (vocaloid):

2/4 = 1/2...

Vocaloid (vocaloid):

two divided by 4 is one half (1/2)

OpenStudy (anonymous):

ohhh so its 1/2? @vocaloid

OpenStudy (anonymous):

or wait -1/2! right? @vocaloid

Vocaloid (vocaloid):

divide the equation by 4 first, and tell me what you get

OpenStudy (anonymous):

1/2x-1y @vocaloid

Vocaloid (vocaloid):

don't forget the 4 at the end

OpenStudy (anonymous):

0 @vocaloid

Vocaloid (vocaloid):

what is 4/4?

OpenStudy (anonymous):

1

Vocaloid (vocaloid):

good, so (1/2)x - y = 1 now add y to each side

OpenStudy (anonymous):

1/2x=2... @vocaloid

OpenStudy (anonymous):

?

Vocaloid (vocaloid):

not quite, remember that you cannnot add 1 + y to get 2 (1/2)x = y + 1 so, what is the slope?

OpenStudy (anonymous):

1/2... @vocaloid

OpenStudy (anonymous):

?

Vocaloid (vocaloid):

yes